# Theoretical Foundations of Prompt Engineering: From Heuristics to Expressivity

### v1

@@ 7.2 Execution: Layerwise Emulation
for every ReLU MLP $N$ satisfying Definition 7.1, there exists a prompt $p_{N}=Enc(N)\in\mathbb{R}^{L_{\max}\times d_{p}}$ (using zero-padding if needed) for which $\sup_{x\in X}\|F_{\theta^{\ast}}(p_{N},x)-N(x)\|\leq\eta$.
@@
We implement the forward pass of $N$ via macro-steps for layers $\ell=1,\dots,S-1$. At macro-step $\ell$, the work token forms queries so that, by the key design $\kappa$, the correct slot attains an $(i^{\star},\Delta)$-margin. Lemma 6.2 then yields exponentially accurate retrieval, and it suffices to pick $\tau$ so that $(L-1)e^{-\Delta/\tau}\leq\eta_{\mathrm{route}}$.
On the resulting compact set of register states and retrieved fragments (from bounded parameters and compact $\mathcal{X}$), Lemma 6.6 gives a token-wise FFN approximating the intended local update map $G_{\ell}$ within the arithmetic budget. Finally, Lemma 6.8 allocates per-step budgets so that routing and arithmetic errors accumulate to at most $\eta$, yielding $\sup_{x\in\mathcal{X}}\lvert F_{\theta^{\ast}}(p_{N},x)-N(x)\rvert\leq\eta$. ∎
@@ Appendix B Simplified Transformer Architecture
We take the prompt width to match the model width, i.e., $d_{p}=D$, so a prompt is a length-$L$ token sequence $p=(p_{1},\ldots,p_{L})\in P=\mathbb{R}^{L\times D}$ with $p_{i}\in\mathbb{R}^{D}$. An input $x\in\mathcal{X}\subset\mathbb{R}^{d}$ is embedded into a single input token
@@
$\mathcal{R}_{\tau}(q;P)\;:=\;S_{\tau}(qK_{P}^{\top})\,V_{P}\in\mathbb{R}^{D}.$
In later constructions, we partition the prompt slots into parameter storage and work memory:
$L]\;=\;I_{\mathrm{par}}\,\dot{\cup}\,I_{\mathrm{wrk}}.$
@@ Appendix C Routing Lemma: Prompt-Memory Retrieval
$\mathcal{R}_{\tau}(q;P)\;:=\;S_{\tau}(qK_{P}^{\top})\,V_{P}\in\mathbb{R}^{D},$
@@ E.3 A one-step bound
Assume (7), (8), and (9). Then for any $t$ and any $(p,x)\in P\times X$,
$\|\widehat{s}_{t+1}-s_{t+1}\|_{\infty}\;\leq\;\|\widehat{s}_{t}-s_{t}\|_{\infty}\;+\;\delta^{\mathrm{arith}}_{t}\;+\;L_{t}\,\delta^{\mathrm{route}}_{t}.$
@@
The middle term is bounded by $\delta^{\mathrm{arith}}_{t}$ by (8). The first term is bounded by $L_{t}\|\widehat{\lambda}_{t}-\lambda_{t}\|_{\infty}\leq L_{t}\delta^{\mathrm{route}}_{t}$ using (9) (absorbing the realized-vs-ideal dependence into the same sensitivity bound). The last term is bounded by $\|\widehat{s}_{t}-s_{t}\|_{\infty}$ by 1-Lipschitzness of the state update in the work-state argument; in our constructions the work-state enters through residual additions and coordinate-wise bounded operations, and this bound can be enforced by scaling if needed. Combining yields the claim. ∎
@@
If $\Psi_{t}$ is not 1-Lipschitz in $s$, replace the last term by $K_{t}\|\widehat{s}_{t}-s_{t}\|_{\infty}$ with the appropriate constant $K_{t}$, and the multi-step bound below becomes a product/sum bound. In our executor design we ensure $K_{t}\leq 1$ by scaling and residual structure, to keep bookkeeping additive.
@@ F.2 Address labels and a fixed key codebook
Let $\mathcal{L}$ be a finite set of address labels indexing the records that may be retrieved by the executor (e.g., ”row $i$ of a matrix used at macro-step tt”, ”bias term”, etc.). We fix a codebook of key vectors $\{k(\ell)\in\mathbb{R}^{D}:\ell\in\mathcal{L}\}$ such that there exists a separation margin $\gamma\in(0,1]$ with
@@
For finite $\mathcal{L}$, such a codebook exists whenever $D$ is sufficiently large (e.g., using nearly-orthogonal vectors). The executor will generate a query $q(\ell)$ that aligns with $k(\ell)$, so that the score margin required by the routing lemma (Appendix C) holds.
@@
A particularly simple choice is to take $\mathcal{L}\subseteq[D]$ and set $k(\ell)=e_{\ell}$, in which case (11) holds with $\gamma=1$.
@@ Appendix G Executor Program: Macro-step Layout and Invariants
This appendix describes the executor as a finite ”program” executed over a sequence of macro-steps. The goal is to (i) make explicit which subroutines are performed at each step (routing reads, arithmetic updates, optional write-backs), and (ii) state a small set of invariants that are maintained throughout execution. These invariants serve as the bookkeeping interface between the macro-step description and the error analysis: they specify what must be true at each step so that routing and token-wise computation can be composed with a controlled end-to-end error.
@@ G.2 Macro-step layout
We organize the depth-$T$ Transformer into $T_{\mathrm{macro}}$ macro-steps. Each macro-step consists of a fixed pattern:
@@
    (Query formation) From the current work token, an affine (or simple FFN) subroutine writes a query vector $q_{t,r}$ into the query register for read $r\in\{1,\ldots,R_{t}\}$.
@@
    where $\mathrm{Inject}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is a fixed linear map selecting the coordinates/registers that receive the read vector. (This is implemented by $W_{V}$ and subsequent fixed linear maps.)
@@ G.3 Execution invariants
We state the invariants that are maintained for all macro-steps $t\in\{0,\ldots,T_{\mathrm{macro}}\}$.
Inv-1 (Boundedness).
All designated token vectors remain in a bounded box: there exists $B>0$ such that
@@
Each subroutine only writes to intended registers: query formation only changes $\Pi_{\mathrm{qry}}$, routing injection only affects designated payload registers, and the arithmetic FFN updates $\Pi_{\mathrm{st}}$ according to the intended $\Psi_{t}$ while keeping all registers bounded (consistent with Inv-1).
G.4 Maintaining invariants and deriving per-step bounds
The next lemma shows that Inv-3 implies a uniform bound on each routed read, and that Inv-1 ensures the arithmetic lemma applies at every step.
@@ G.4 Maintaining invariants and deriving per-step bounds
Assume Inv-1–Inv-3 hold at macro-step $t$. Let $q_{t,r}$ be a query with margin $\Delta_{t,r}$ for its target slot, and assume all prompt values satisfy $\|(V_{P})_{j,:}\|_{\infty}\leq B$. Then the routed read obeys
@@
The routing bound is exactly Lemma C.2 (Appendix C). The arithmetic prerequisite follows since Inv-1 ensures the FFN inputs remain in $\mathcal{B}_{B}$, the compact domain on which Appendix D provides uniform approximation. ∎
G.5 A generic macro-step error statement
@@ G.6 End-to-end view
$\sup_{x\in X}\bigl|F_{\theta^{\ast}}(p,x)-\widetilde{F}_{\theta^{\ast}}(p,x)\bigr|\;\leq\;\varepsilon_{\mathrm{exec}},$
where $\widetilde{F}_{\theta^{\ast}}$ denotes the idealized program output under perfect routing and arithmetic. This is the executor-side half of the triangle-inequality split used in the main proof.
@@
To fully instantiate this appendix for the final proof, the main construction must specify: (i) the label set $\{\ell_{t,r}\}$ and memory layout in $\mathcal{I}_{\mathrm{par}}$, (ii) how queries $q_{t,r}$ are generated and why they satisfy the margin condition (Inv-3), and (iii) why boundedness (Inv-1) holds under the chosen encodings and updates. These are precisely the points addressed by the routing, encoding, arithmetic, and composition appendices.
@@ Appendix H Full Proof of the Main Theorem
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be compact as in Assumption 3.1. There exists a fixed set of executor parameters $\theta^{\ast}$ (a simplified Transformer as in Appendix B) such that for every $g\in C(X)$ and every $\varepsilon>0$ there exists a soft prompt $p_{g}\in\mathcal{P}$ with
$\sup_{x\in X}\bigl|F_{\theta^{\ast}}(p_{g},x)-g(x)\bigr|\;<\;\varepsilon.$
Equivalently, for the prompt-induced family $H_{\theta^{\ast}}=\{F_{\theta^{\ast}}(p,\cdot):p\in P\}$ (Definition 3.4), we have $\overline{H_{\theta^{\ast}}}^{\|\cdot\|_{\infty}}=C(X)$.
@@
Fix an arbitrary target $g\in C(X)$ and $\varepsilon>0$. Set
@@
By the classical universal approximation theorem for ReLU networks on compact domains, there exists a (finite) ReLU MLP $N_{g}:X\to\mathbb{R}$ such that
$\sup_{x\in X}|N_{g}(x)-g(x)|\;\leq\;\varepsilon_{\mathrm{approx}}.$ (13)
Without loss of generality, choose $N_{g}$ to be a depth-$S=3$ (one-hidden-layer) ReLU MLP as in Definition 7.1, which is sufficient by the classical UAT on compact $X$. Write its parameters in the same format as Definition 7.1 and collect them into a finite list
$\pi_{g}:=\big(W^{(1)},b^{(1)},a,c\big),$
@@
Each record label specifies which fragment is needed at which macro-step (e.g., ”layer $\ell$ weight row ii”, ”layer $\ell$ bias block”), and each payload is produced by a fixed packing map. This yields a prompt matrix $P_{g}\in\mathbb{R}^{L\times D}$ and hence a prompt $p_{g}\in P$.
@@
$\delta^{\mathrm{arith}}_{t}\equiv\frac{\varepsilon_{\mathrm{exec}}}{4T_{\mathrm{macro}}},\qquad L_{t}\,\delta^{\mathrm{route}}_{t}\equiv\frac{\varepsilon_{\mathrm{exec}}}{4T_{\mathrm{macro}}},\qquad t=0,\ldots,T_{\mathrm{macro}}-1,$ (15)
@@
Routing. For each read in each macro-step, the executor forms a query $q$ targeting the corresponding stored label. By the margin invariant in Appendix G and the routing bound in Appendix C, we can choose a temperature $\tau$ (and, if needed, scaling inside the fixed query generation) so that every read has impurity at most $\rho$, hence each read error is at most $2B\rho$ in $\|\cdot\|_{\infty}$ for a uniform value bound $B$. Since each macro-step uses at most $R$ reads and the subsequent assembly is a fixed linear map, this yields a bound of the form
$\|\widehat{\lambda}_{t}(p_{g},x)-\lambda_{t}(p_{g})\|_{\infty}\;\leq\;\delta^{\mathrm{route}}_{t}\qquad\text{uniformly over }x\in X,$ (16)
by choosing $\rho$ (equivalently $\tau$) small enough. (A fully explicit inequality is given by the routing lemma and its corollaries.)
@@
$\sup_{x\in X}\bigl\|\widehat{\Psi}_{t}(\widehat{s}_{t},\lambda_{t}(p_{g}))-\Psi_{t}(\widehat{s}_{t},\lambda_{t}(p_{g}))\bigr\|_{\infty}\;\leq\;\delta^{\mathrm{arith}}_{t}.$ (17)
@@
$\sup_{x\in X}\ \|\widehat{s}_{T_{\mathrm{macro}}}(x)-s_{T_{\mathrm{macro}}}(x)\|_{\infty}\;\leq\;\sum_{t=0}^{T_{\mathrm{macro}}-1}\left(\delta^{\mathrm{arith}}_{t}+L_{t}\,\delta^{\mathrm{route}}_{t}\right)\;\leq\;\varepsilon_{\mathrm{exec}}/2.$ (18)
@@
$\sup_{x\in X}\bigl|F_{\theta^{\ast}}(p_{g},x)-N_{g}(x)\bigr|\;\leq\;\varepsilon_{\mathrm{exec}}.$ (19)
@@
$\sup_{x\in X}\bigl|F_{\theta^{\ast}}(p_{g},x)-g(x)\bigr|\;\leq\;\sup_{x\in X}\bigl|F_{\theta^{\ast}}(p_{g},x)-N_{g}(x)\bigr|+\sup_{x\in X}\bigl|N_{g}(x)-g(x)\bigr|\;\leq\;\varepsilon_{\mathrm{exec}}+\varepsilon_{\mathrm{approx}}\;=\;\varepsilon.$
Since $g\in C(X)$ and $\varepsilon>0$ were arbitrary, the theorem follows. ∎

### v2

@@ 7.2 Execution: Layerwise Emulation
for every ReLU MLP $N$ satisfying Definition 7.1, there exists a prompt $p_{N}=\mathrm{Enc}(N)\in\mathbb{R}^{L_{\max}\times D}$ (using zero-padding if needed) such that
$\sup_{x\in\mathcal{X}}\big|F_{\theta^{\ast}}(p_{N},x)-N(x)\big|\leq\eta.$
@@
We implement the forward pass of $N$ via macro-steps for layers $\ell=1,\dots,S-1$. At macro-step $\ell$, the work token forms queries so that, by the key design $\kappa$, the correct slot attains an $(i^{\star},\Delta)$-margin. Lemma 6.2 then yields exponentially accurate retrieval, and it suffices to pick $\tau$ so that $(L_{\max}-1)e^{-\Delta/\tau}\leq\eta_{\mathrm{route}}$.
On the resulting compact set of register states and retrieved fragments (from bounded parameters and compact $\mathcal{X}$), Lemma 6.6 gives a token-wise FFN approximating the intended local update map $G_{\ell}$ within the arithmetic budget. Finally, Lemma 6.8 allocates per-step budgets so that routing and arithmetic errors accumulate to at most $\eta$, yielding $\sup_{x\in\mathcal{X}}\big|F_{\theta^{\ast}}(p_{N},x)-N(x)\big|\leq\eta$. ∎
@@ Appendix B Simplified Transformer Architecture
We take the prompt width to match the model width, i.e., $d_{p}=D$, so a prompt is a length-$L$ token sequence $p=(p_{1},\ldots,p_{L})\in\mathcal{P}=\mathbb{R}^{L\times D}$ with $p_{i}\in\mathbb{R}^{D}$. An input $x\in\mathcal{X}\subset\mathbb{R}^{d}$ is embedded into a single input token
@@
$\mathcal{R}_{\tau}(q;P)\;:=\;S_{\tau}\!\left(\frac{qK_{P}^{\top}}{\sqrt{D}}\right)\,V_{P}\in\mathbb{R}^{D}.$
(As in self-attention, the $\sqrt{D}$ factor can be absorbed into $\tau$; we keep it explicit for consistency with $A_{\tau}$.)
@@ Appendix C Routing Lemma: Prompt-Memory Retrieval
$\mathcal{R}_{\tau}(q;P)\;:=\;S_{\tau}\!\left(\frac{qK_{P}^{\top}}{\sqrt{D}}\right)\,V_{P}\in\mathbb{R}^{D},$
@@ E.3 A one-step bound
Assume (7), (8), and (9). In addition, assume that the ideal update $\Psi_{t}$ is $K_{t}$-Lipschitz in the work-state argument on the relevant domain:
$\|\Psi_{t}(s,\lambda)-\Psi_{t}(s^{\prime},\lambda)\|_{\infty}\leq K_{t}\|s-s^{\prime}\|_{\infty}\qquad\text{for all }s,s^{\prime}\text{ and admissible }\lambda.$ (10)
Then for any $t$ and any $(p,x)\in P\times X$,
$\|\widehat{s}_{t+1}-s_{t+1}\|_{\infty}\;\leq\;K_{t}\,\|\widehat{s}_{t}-s_{t}\|_{\infty}\;+\;\delta^{\mathrm{arith}}_{t}\;+\;L_{t}\,\delta^{\mathrm{route}}_{t}.$
In particular, if $K_{t}\leq 1$ (as enforced in our executor design), this reduces to the additive bound $\|\widehat{s}_{t+1}-s_{t+1}\|_{\infty}\leq\|\widehat{s}_{t}-s_{t}\|_{\infty}+\delta^{\mathrm{arith}}_{t}+L_{t}\delta^{\mathrm{route}}_{t}$.
@@
The middle term is bounded by $\delta^{\mathrm{arith}}_{t}$ by (8). The first term is bounded by $L_{t}\|\widehat{\lambda}_{t}-\lambda_{t}\|_{\infty}\leq L_{t}\delta^{\mathrm{route}}_{t}$ using (9) (absorbing the realized-vs-ideal dependence into the same sensitivity bound). The last term is bounded by $K_{t}\|\widehat{s}_{t}-s_{t}\|_{\infty}$ by (10). Combining yields the claim. ∎
@@
The constant $K_{t}$ in (10) depends on the ideal update $\Psi_{t}$ and the domain. If one prefers to keep the bookkeeping additive, it suffices to enforce $K_{t}\leq 1$. In our executor design this can be achieved (when needed) by scaling and residual structure, since the work-state typically enters through residual additions and coordinate-wise bounded operations.
@@ F.2 Address labels and a fixed key codebook
Let $\mathcal{L}$ be a finite set of address labels indexing the records that may be retrieved by the executor (e.g., ”row $i$ of a matrix used at macro-step tt”, ”bias term”, etc.). Fix a key dimension $d_{k}\leq D$ and a codebook of key vectors $\{\kappa(\ell)\in\mathbb{R}^{d_{k}}:\ell\in\mathcal{L}\}$ such that there exists a separation margin $\gamma\in(0,1]$ with
@@
For finite $\mathcal{L}$, such a codebook exists whenever $d_{k}$ is sufficiently large (e.g., using nearly-orthogonal vectors). The executor will generate a query $q(\ell)$ whose key block aligns with $\kappa(\ell)$, so that the score margin required by the routing lemma (Appendix C) holds.
@@
A particularly simple choice is to take $\mathcal{L}\subseteq[d_{k}]$ and set $\kappa(\ell)=e_{\ell}$, in which case (12) holds with $\gamma=1$.
@@ Appendix G Executor Program: Macro-step Layout and Invariants
This appendix describes the executor as a finite “program” executed over a sequence of macro-steps. Our goals are: (i) to make explicit which subroutines are performed at each step (routing reads, arithmetic updates, optional write-backs), and (ii) to state a small set of invariants maintained throughout execution. These invariants act as the bookkeeping interface between the macro-step description and the error analysis: they specify what must hold at each step so that routing and token-wise computation can be composed with a controlled end-to-end error.
Admissible prompts (uniformity domain).
Throughout Appendix G, we restrict attention to the admissible prompt set $\mathcal{P}_{\mathrm{adm}}\subseteq\mathbb{R}^{L\times D}$ produced by the encoding scheme of Section 4.2 and Appendix F. In particular, prompts in $\mathcal{P}_{\mathrm{adm}}$ store records in fixed bounded ranges determined by the construction. All invariants and bounds in this appendix are uniform over $\mathcal{P}_{\mathrm{adm}}$ and $x\in\mathcal{X}$. See Remark 3.5 for the relation between the ambient prompt space $\mathcal{P}$ and the admissible subset $\mathcal{P}_{\mathrm{adm}}$ used for uniform bounds.
@@ G.2 Macro-step layout
We organize the depth-$T$ Transformer into $T_{\mathrm{macro}}$ macro-steps. Each macro-step consists of the following fixed pattern of subroutines:
@@
    (Query formation) From the current work token, a fixed affine (or simple FFN) subroutine writes a query vector $q_{t,r}$ into the query register for read $r\in\{1,\ldots,R_{t}\}$.
@@
    where $\mathrm{Inject}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is a fixed linear map selecting the coordinates/registers that receive the read vector (implemented by $W_{V}$ and subsequent fixed linear maps).
@@ G.3 Execution invariants
We now state the invariants maintained for all macro-steps $t\in\{0,\ldots,T_{\mathrm{macro}}\}$. All statements below are uniform over $P\in\mathcal{P}_{\mathrm{adm}}$ and $x\in\mathcal{X}$.
Inv-1 (Uniform boundedness).
There exists a bound $B>0$ depending only on the fixed executor architecture and construction constants (in particular on $\theta^{\ast}$, $L,D,T_{\mathrm{macro}}$, and the encoding ranges defining $\mathcal{P}_{\mathrm{adm}}$), such that all designated token vectors remain in the box $\mathcal{B}_{B}=[-B,B]^{D}$:
@@
In particular, prompt tokens satisfy $\|p_{j}\|_{\infty}\leq B$ for all $j\in[L]$, and since the value map $W_{V}$ is fixed in $\theta^{\ast}$, the prompt values $(V_{P})_{j,:}$ are also uniformly bounded (up to a constant factor absorbed into $B$), as required by the routing bound.
@@
Each subroutine only writes to intended registers: query formation only changes $\Pi_{\mathrm{qry}}$, routing injection only affects designated payload registers, and the arithmetic FFN updates $\Pi_{\mathrm{st}}$ according to the intended $\Psi_{t}$ while preserving boundedness (consistent with Inv-1).
G.4 Maintaining Inv-1: a generic sufficient condition
The next lemma records a clean sufficient condition under which the uniform boundedness invariant (Inv-1) holds. Its role is to align the boundedness bookkeeping with the actual executor dynamics (which includes residual additions).
Lemma G.1 (Sufficient condition for Inv-1).
Assume that the construction fixes $B_{0}>0$ such that:
  1. (i)
    The encoding maps used to form the initial tokens $(P,u(x),w,o)$ produce vectors in $\mathcal{B}_{B_{0}}$ uniformly over $x\in\mathcal{X}$ and all $P\in\mathcal{P}_{\mathrm{adm}}$.
  2. (ii)
    Each macro-step subroutine (query formation, routing injection, linear assembly, and arithmetic FFN) is implemented by fixed weights chosen so that $\mathcal{B}_{B_{0}}$ is an invariant set for the executor dynamics, i.e., each subroutine maps inputs in $\mathcal{B}_{B_{0}}$ to outputs in $\mathcal{B}_{B_{0}}$. (For instance, this can be ensured by range-restricted encodings together with appropriate scaling of linear maps so that residual additions remain within $\mathcal{B}_{B_{0}}$ on that set; explicit clipping/saturation is one sufficient but not necessary mechanism.)
Then Inv-1 holds with $B=B_{0}$.
Proof.
By induction over macro-steps. The base case follows from (i). For the inductive step, assume all designated tokens lie in $\mathcal{B}_{B_{0}}$ at macro-time $t$. Applying the subroutines of macro-step $t$ preserves membership in $\mathcal{B}_{B_{0}}$ by (ii), hence the designated tokens remain bounded at macro-time $t+1$. ∎
G.5 Maintaining invariants and deriving per-step bounds
The next lemma shows that Inv-3 implies an exponentially small routing error for each read, and that Inv-1 guarantees that the arithmetic approximation can be taken uniformly on a fixed compact domain.
@@ G.5 Maintaining invariants and deriving per-step bounds
Assume Inv-1–Inv-3 hold at macro-step $t$. Let $q_{t,r}$ be a query with margin $\Delta_{t,r}$ for its target slot. Then the routed read obeys
@@
The routing bound is exactly Lemma C.2 (Appendix C). The arithmetic prerequisite follows since Inv-1 ensures that all FFN inputs remain in $\mathcal{B}_{B}$, the compact domain on which Appendix D provides uniform approximation. ∎
G.6 A generic macro-step error statement
@@ G.7 End-to-end view
$\sup_{x\in\mathcal{X}}\bigl|F_{\theta^{\ast}}(p,x)-\widetilde{F}_{\theta^{\ast}}(p,x)\bigr|\;\leq\;\varepsilon_{\mathrm{exec}},$
where $\widetilde{F}_{\theta^{\ast}}$ denotes the idealized program output under perfect routing and perfect arithmetic. This is the executor-side half of the triangle-inequality split used in the main proof.
@@
To fully instantiate this appendix for the final proof, the main construction must specify: (i) the label set $\{\ell_{t,r}\}$ and memory layout in $\mathcal{I}_{\mathrm{par}}$, (ii) how queries $q_{t,r}$ are generated and why they satisfy the margin condition (Inv-3), and (iii) why the uniform boundedness (Inv-1) holds with a constant $B$ uniform over all admissible prompts in $\mathcal{P}_{\mathrm{adm}}$, under the chosen encodings, injections, and arithmetic updates. These are precisely the points addressed by the routing, encoding, arithmetic, and composition appendices.
@@ Appendix H Full Proof of the Main Theorem
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be compact as in Assumption 3.1. There exists a fixed set of executor parameters $\theta^{\ast}$ (a simplified Transformer as in Appendix B) such that for every $g\in C(\mathcal{X})$ and every $\varepsilon>0$ there exists a soft prompt $p_{g}\in\mathcal{P}$ with
$\sup_{x\in\mathcal{X}}\bigl|F_{\theta^{\ast}}(p_{g},x)-g(x)\bigr|\;<\;\varepsilon.$
Equivalently, for the prompt-induced family $H_{\theta^{\ast}}=\{F_{\theta^{\ast}}(p,\cdot):p\in\mathcal{P}\}$ (Definition 3.4), we have $\overline{H_{\theta^{\ast}}}^{\|\cdot\|_{\infty}}=C(\mathcal{X})$.
@@
Fix an arbitrary target $g\in C(\mathcal{X})$ and $\varepsilon>0$. Set
@@
By the classical universal approximation theorem for ReLU networks on compact domains, there exists a (finite) ReLU MLP $N_{g}:\mathcal{X}\to\mathbb{R}$ such that
$\sup_{x\in\mathcal{X}}|N_{g}(x)-g(x)|\;\leq\;\varepsilon_{\mathrm{approx}}.$ (14)
Without loss of generality, choose $N_{g}$ to be a depth-$S=3$ (one-hidden-layer) ReLU MLP in the sense of Definition 7.1. In particular, write
$N_{g}(x)=a^{\top}\sigma(Wx+b)+c\quad\text{and identify}\quad W^{(1)}=W,\ b^{(1)}=b,\ W^{(2)}=a^{\top},\ b^{(2)}=c.$
Collect the parameters in the same format as Definition 7.1:
$\pi_{g}:=\big(W^{(1)},b^{(1)},W^{(2)},b^{(2)}\big),$
@@
Each record label specifies which fragment is needed at which macro-step (e.g., ”layer $\ell$ weight row ii”, ”layer $\ell$ bias block”), and each payload is produced by a fixed packing map. This yields a prompt matrix $p_{g}\in\mathbb{R}^{L\times D}$ (equivalently, $p_{g}\in\mathcal{P}$).
@@
$\delta^{\mathrm{arith}}_{t}\equiv\frac{\varepsilon_{\mathrm{exec}}}{2T_{\mathrm{macro}}},\qquad L_{t}\,\delta^{\mathrm{route}}_{t}\equiv\frac{\varepsilon_{\mathrm{exec}}}{2T_{\mathrm{macro}}},\qquad t=0,\ldots,T_{\mathrm{macro}}-1,$ (16)
@@
Routing. For each read in each macro-step, the executor forms a query $q$ targeting the corresponding stored label. By the margin invariant in Appendix G and the routing bound in Appendix C, we can choose a temperature $\tau$ (and, if needed, scaling inside the fixed query generation) so that every read has impurity at most $\rho$, hence each read error is at most $2B\rho$ in $\|\cdot\|_{\infty}$ for a uniform value bound $B$. Since each macro-step uses at most $R$ reads and the subsequent assembly is a fixed linear map, we can choose $\rho$ (equivalently $\tau$) small enough so that the assembled parameter error satisfies
$\|\widehat{\lambda}_{t}(p_{g},x)-\lambda_{t}(p_{g})\|_{\infty}\;\leq\;\delta^{\mathrm{route}}_{t}\qquad\text{uniformly over }x\in\mathcal{X}.$ (17)
@@
$\sup_{x\in\mathcal{X}}\bigl\|\widehat{\Psi}_{t}(\widehat{s}_{t},\lambda_{t}(p_{g}))-\Psi_{t}(\widehat{s}_{t},\lambda_{t}(p_{g}))\bigr\|_{\infty}\;\leq\;\delta^{\mathrm{arith}}_{t}.$ (18)
@@
$\sup_{x\in\mathcal{X}}\ \|\widehat{s}_{T_{\mathrm{macro}}}(x)-s_{T_{\mathrm{macro}}}(x)\|_{\infty}\;\leq\;\sum_{t=0}^{T_{\mathrm{macro}}-1}\left(\delta^{\mathrm{arith}}_{t}+L_{t}\,\delta^{\mathrm{route}}_{t}\right)\;\leq\;\varepsilon_{\mathrm{exec}}.$ (19)
@@
$\sup_{x\in\mathcal{X}}\bigl|F_{\theta^{\ast}}(p_{g},x)-N_{g}(x)\bigr|\;\leq\;\varepsilon_{\mathrm{exec}}.$ (20)
@@
$\sup_{x\in\mathcal{X}}\bigl|F_{\theta^{\ast}}(p_{g},x)-g(x)\bigr|\;\leq\;\sup_{x\in\mathcal{X}}\bigl|F_{\theta^{\ast}}(p_{g},x)-N_{g}(x)\bigr|+\sup_{x\in\mathcal{X}}\bigl|N_{g}(x)-g(x)\bigr|\;\leq\;\varepsilon_{\mathrm{exec}}+\varepsilon_{\mathrm{approx}}\;=\;\varepsilon.$
Since $g\in C(\mathcal{X})$ and $\varepsilon>0$ were arbitrary, the theorem follows. ∎